# CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light Curves

### v1

@@ 4.9.2 Feature of a Curve - II
Clustering Method Cluster Label TOIs Curves Percentage (%)
@@
4.9.3 Comparison and Interpretation of Clustering Outcomes
To explore the morphological diversity among the 620 planet-like candidates, we applied a range of clustering algorithms to a 10-dimensional feature set engineered from phase-folded light curves. This feature set is as described in section 4.9.2. Our objective was to determine whether these features could reveal structural subgroups and isolate known TOIs within well-defined clusters. Results of clustering methods are shown in Table 7
Among all tested methods, Gaussian Mixture Models (GMM) with two components produced the clearest separation: one cluster contained 17 TOIs among 155 curves, achieving the highest TOI density. Dirichlet Process GMM (DPMM) performed comparably well, grouping 16 TOIs within a smaller cluster of 114 curves, indicating a stronger enrichment but with more fragmentation overall.
While other methods such as DBSCAN, HDBSCAN, and KMeans provided partial clustering success, they either included fewer TOIs or distributed them across broader clusters. The probabilistic nature of GMM and DPMM allowed them to better model the subtle variations in transit morphology, making them more suitable for this task.
In conclusion, the combination of interpretable features and flexible model-based clustering yielded the best separation of high-quality transit-like signals. The GMM and DPMM results (sampled in figures 34, 31, 32, and 33 to visualize morphological similarity) represent the most promising configurations to guide further candidate filtering, validation, and prioritization.
@@ 4.9.3 Comparison and Interpretation of Clustering Outcomes
Figure 33: Random sample from gmm cluster 1
Figure 34: Random sample from dpmm cluster 0
5 Computational Resources and Runtime
@@ 6.1 Synthesis of Key Findings
The CLARA framework demonstrates that synthetic set design can systematically control the behavior of Unsupervised Random Forests for astronomical anomaly detection, addressing two fundamental challenges in large-scale time-series surveys: the need for interpretable, steerable and computationally efficient discovery mechanisms and the validation of detected anomalies against genuine astrophysical phenomena.
Our investigation reveals several critical insights into the relationship between synthetic training parameters and model performance. The systematic parameter exploration across 36 URF-4 subvariants establishes that given our six performance metrics, transit duration and noise level emerge as the most influential factors for steering model behavior, with transit duration dominating both TOI Recall and Anomaly Rate metrics while noise level primarily affects importance-based prioritization. This finding has immediate practical implications: researchers can now predictably tune anomaly detection systems toward specific scientific objectives by adjusting these key parameters within viable operational ranges (50-300 ppm for noise, 13-27 days for duration). More finer mapping of feature to performance metric using the a suitable combined scoring metric using a single $\alpha$ coefficient or multiple coefficients can be done to predict urf behavior better.
The $\alpha$-parameterized scoring framework successfully enables targeted control over the scientific trade-off between detection completeness and astrophysical significance. Cross-sector validation across TESS SPOC Sectors 1-5 confirms that this behavioral steering mechanism generalizes robustly, with $\alpha$=0.9 models consistently maximizing binary TOI recall (64% approx) and $\alpha$=0.3 models concentrating high-value targets within narrow score ranges (1.6-9.7% percentile span (approx)). This consistency across diverse stellar populations validates the transferability of synthetic training approaches and enables confident deployment without sector-specific retraining, especially when a combined anomaly scoring metric like N-WRSS is used.
@@ 6.2 Astrophysical Validation and Physical Interpretability
The correlation analysis between anomaly scores and stellar parameters provides compelling evidence that URF-detected anomalies correspond to genuine astrophysical phenomena rather than instrumental artifacts. The t-SNE embedding analysis using astrophysical features [Teff, vtan, ruwe] reveals that high-scoring anomalies cluster along distinct manifolds in parameter space, suggesting they represent a physically coherent subpopulation rather than random noise.
Particularly significant is the morphological classification validation through SIMBAD cross-matching and cosine similarity analysis. The complete absence of high-anomaly candidates (N-WRSS $\geq$ 0.85) with low morphological similarity (cosine similarity $\leq$ 0.3) to known astrophysical classes strongly supports the hypothesis that URF anomaly scores reflect genuine astrophysical structure. This finding addresses a fundamental concern in unsupervised anomaly detection: whether statistical "weirdness" corresponds to scientifically meaningful phenomena. The consistent morphological similarity of high-scoring candidates to established planetary, stellar, and binary star classes validates the physical realism of the detection framework.
The Normalized Weighted Root Sum of Squares (N-WRSS) scoring scheme effectively combines the complementary strengths of conservative ($\alpha$=0.3) and aggressive ($\alpha$=0.9) models, providing a unified metric that captures both consensus-based confidence and comprehensive coverage. The distribution analysis across SIMBAD object classes reveals that planet-like candidates maintain higher representation at elevated N-WRSS thresholds compared to stellar or binary classifications, suggesting that the framework exhibits appropriate sensitivity to transit-like morphologies.
6.3 Computational Efficiency and Methodological Advances
The CLARA framework introduces several methodological innovations that advance the state of unsupervised astronomical anomaly detection. The systematic exploration of synthetic parameter space moves beyond the fixed hyperparameter approaches of previous work, enabling researchers to understand and control the mapping between input design choices and detection outcomes. This controllability represents a paradigm shift from black-box anomaly detection toward interpretable, goal-oriented discovery systems.
The computational efficiency achieved on modest hardware (4-core i3 CPU, 32GB RAM) demonstrates the practical scalability of the approach. Processing 1̃6,000 to 20,000 light curves per sector in approximately 80 minutes to 120 minutes per $\alpha$ model (it varies on number of curves and value of $\alpha$ - higher alpha models take lesser time than lower alpha models which behave more conservatively), with full anomaly scoring, feature extraction, and morphological classification completing within reasonable timeframes, validates the framework’s suitability for large-scale surveys.
The modular pipeline design facilitates extensibility to other survey architectures and scientific objectives. The clear separation between synthetic set generation, anomaly scoring, and astrophysical validation enables researchers to adapt individual components while preserving the overall framework integrity. This modularity is particularly valuable for extending CLARA to Full Frame Image (FFI) data, where the broader stellar population and different noise characteristics would require adjusted synthetic training strategies while maintaining the core methodological approach.
@@ 6.4 Limitations and Scope of Applicability
Several important limitations constrain the generalization of current results. The restriction to 2-minute cadence SPOC targets introduces selection bias toward brighter, pre-selected stars, potentially limiting applicability to the broader TESS population accessible through FFI data. Our analysis suggests that FFI-based light curves would establish different baselines for "normal" behavior, requiring recalibration of anomaly score thresholds and potentially altering the specific mappings between synthetic parameters and performance metrics. While the core hypothesis, that synthetic set design steers model behavior, should remain valid, the quantitative relationships established here may not transfer directly to FFI populations.
The evaluation framework relies heavily on TOIs as ground-truth proxies, which introduces potential circularity in performance assessment. TOIs themselves represent a biased sample, selected through existing detection algorithms and subject to observational and instrumental selection effects. This dependence limits our ability to assess the discovery of genuinely novel phenomena that fall outside current detection paradigms. The morphological classification via SIMBAD cross-matching partially addresses this limitation by providing independent astrophysical validation, but the sparse labeling in SIMBAD restricts the scope of this validation approach.
The synthetic light curve generation employs simplified transit models (box-shaped and Mandel-Agol) with uniform parameter distributions, which may not adequately represent the full diversity of real planetary systems. More sophisticated synthetic generation incorporating realistic parameter correlations, stellar activity, and instrumental systematics, if possible to be tested under controlled changes in the parameter values, could improve model performance and generalization. Similarly, the feature extraction approach, while computationally efficient, truncates light curves to 3000 flux points and may miss long-duration or late-occurring anomalies.
@@ 6.5 Future Directions and Extensions
The CLARA framework establishes a foundation for several promising research directions. Based on the comprehensive table of planned improvements (see author’s development notes), we outline nine key areas for future investigation:
@@
Immediate Technical Extensions: The framework will be scaled to a proposed 500,000 light curve benchmark. This will test the scalability of the CLARA pipeline at clara-cli.
Enhanced Statistical Validation: Future work will implement rigorous statistical validation frameworks to quantify detection significance and establish confidence intervals for anomaly classifications. This includes comparison studies with established methods like Crake & Martínez-Galarza (2023), Melton et al. (2024) etc. incorporating their insights. Additionally, correlation analysis between stellar parameters (Teff, RUWE, vtan) and N-WRSS scores, more specifically, quantifying why the stellar parameters that correlate well with better high N-WRSS clustering tend to do so, will provide deeper astrophysical interpretation of detected anomalies.
Advanced Morphological Analysis: The morphological clustering analysis using engineered features represents an initial step toward automated anomaly taxonomy. Planned expansions include: (1) sophisticated feature combination studies beyond the current 10-dimensional representation as in section 4.9.2, (2) integration of deep learning-based representations for more nuanced morphological classification, and (3) development of improved metrics that surpass current TOI-based and other metrics as in Table 3. The combination of unsupervised detection with supervised morphological classification offers a promising hybrid approach for systematic discovery of new astrophysical phenomena.
Cross-Survey Applications: Extension to other time-series surveys (Kepler, LSST, K2) will test the broader applicability of controllable synthetic training approaches. Each survey presents unique challenges in sampling synthetic set features, and astronomical populations, requiring adapted synthetic generation strategies while preserving core framework principles. The modular design of CLARA facilitates such extensions by isolating survey-specific components from the general methodological approach.
Theoretical Framework Development: A key future priority is to formalize the theoretical foundations of the Normalized Weighted Root Sum Score (N-WRSS), which was introduced as a composite anomaly metric combining outputs from independently trained URF models. In its current formulation, weights are applied to the square roots of raw anomaly scores prior to normalization, primarily to reduce score variance and compress outlier influence in a computationally tractable way. However, this approach remains heuristic. A more rigorous theoretical analysis is needed to justify the use of square-root transformations, evaluate the statistical properties of such weighted combinations, and explore alternative aggregation strategies (e.g., geometric means, entropy-based weights, or learned fusion mechanisms).
In parallel, the revised performance scoring function used for alpha-controlled model selection, defined as a convex combination of normalized recall, anomaly suppression, and interpretability metrics, also warrants formal study. Although the current structure guarantees convexity and boundedness, its empirical weighting scheme (based on an $\alpha$-partitioned group structure) lacks statistical justification. Future work will explore principled approaches for assigning metric weights, sensitivity analysis across parameter regimes, and potential use of Pareto front optimization to balance competing objectives in a non-scalar fashion. Establishing rigorous foundations for both anomaly scoring and performance evaluation would enhance generalization beyond specific URF configurations and improve the interpretability and reliability of CLARA’s outputs across diverse stellar populations.
Automated Optimization: Development of meta-optimization deep learning based techniques for automatically learning optimal synthetic parameters represents a natural evolution beyond the current grid-search approach. Machine learning methods for optimizing synthetic set design based on target performance metrics could eliminate manual synthetic set generation parameter tuning and enable adaptive optimization for evolving scientific priorities. This is particularly valuable for multi-objective optimization scenarios where researchers must balance competing detection goals.
GAN-Driven Post-Filtering: A potential extension of CLARA involves training a post-clustering classifier to distinguish genuine transit-like curves from false positives. To avoid reliance on labeled TOIs, this model could be trained using GAN-generated light curves, where positive samples include synthetic Mandel-Agol transits injected into GAN-simulated stellar variability, and negatives consist of non-transit or eclipsing binary morphologies. The resulting classifier (e.g., CNN or contrastive encoder) would serve as a final, learnable filter after URF-based anomaly detection and morphological similarity based clustering (Cosine similarity, DTW, DBSCAN etc.), improving robustness against spurious signals.
6.6 Broader Context and Significance
The CLARA framework contributes to the broader evolution of astronomical methodology toward intelligent, unsupervised discovery systems capable of handling the data volumes generated by modern surveys. By demonstrating that unsupervised anomaly detection can be made both controllable and interpretable, this work addresses fundamental challenges in the transition from human-expert-driven discovery to machine-assisted exploration of astronomical parameter space.
The validation of detected anomalies through independent astrophysical catalogs establishes a methodology for bridging statistical pattern recognition with physical understanding. This integration is crucial for maintaining scientific rigor while leveraging the pattern detection capabilities of machine learning approaches. The framework provides a model for how automated discovery systems can incorporate domain knowledge and external validation to ensure astrophysical relevance of detected patterns.
The successful demonstration of behavioral steering through synthetic training represents a significant advance in making machine learning systems more transparent and controllable for scientific applications. Rather than treating unsupervised anomaly detection as a black-box process, CLARA enables researchers to understand and predict how methodological choices affect discovery outcomes. This transparency is essential for building confidence in automated unsupervised systems and ensuring that scientific priorities drive technological implementation rather than being constrained by algorithmic limitations.
In conclusion, the CLARA framework establishes both methodological foundations and practical tools for controllable, interpretable anomaly detection in large-scale astronomical surveys. The demonstrated ability to systematically steer detection behavior while maintaining astrophysical validity provides a pathway toward more intelligent, goal-oriented discovery systems capable of adapting to evolving scientific priorities and survey characteristics. As astronomical data volumes continue to grow exponentially, such frameworks will become increasingly essential for realizing the full scientific potential of modern survey missions.
@@ 7 Conclusions
We have presented CLARA, a modular and controllable framework for unsupervised transit detection using TESS light curves. Our two-part investigation demonstrates that synthetic training set design plays a critical role in shaping the behavior and effectiveness of Unsupervised Random Forest (URF) models for anomaly detection.
In Part I, we introduced URF-4, a synthetic-injection-based URF variant with tunable parameters, and performed a systematic parameter sweep across multiple synthetic configurations. By defining $\alpha$-dependent scoring metrics and balancing recall with astrophysical importance, we showed that synthetic training sets can be designed to prioritize either detection completeness ($\alpha=0.9$), astrophysical significance ($\alpha=0.3$), or a balanced trade-off ($\alpha=0.5$). We validated the generalization of these models across TESS Sectors 1–5 and demonstrated that synthetic design enables consistent and predictable performance aligned with distinct scientific goals. Our cross-sector generalization tests involved scoring 384,000 light curves (128,000 per $\alpha$-variant) sampled from five distinct TESS sectors, validating CLARA’s consistency across diverse stellar populations and observational conditions.
In Part II, we assessed the astrophysical relevance of URF-identified anomalies using stellar parameters from Gaia DR3, t-SNE cluster embeddings, and morphological similarity matching with SIMBAD-labeled TOIs. We introduced the Normalized Weighted Root Sum Score (N-WRSS) as a hybrid metric for reconciling high- and low-$\alpha$ models. This analysis revealed that high-scoring anomalies are concentrated in distinct astrophysical subspaces (e.g., in $\mathrm{T_{eff}}$$\mathrm{RUWE}$$\mathrm{v_{tan}}$ space), and exhibit morphological similarity to known exoplanetary and stellar variability classes.
Cosine similarity-based weak labeling showed that a majority of high-scoring anomalies align with planet-like or stellar variability categories, while no high-anomaly, low-similarity candidates were found - indicating the physical plausibility of detected anomalies and minimizing the risk of false positives due to noise.
Together, these results validate CLARA as a physically grounded, interpretable, and generalizable anomaly detection pipeline. By combining synthetic data-driven tuning, astrophysical correlation analysis, and unsupervised clustering techniques, CLARA offers a powerful toolkit for large-scale, label-free exploration of transit-like signals in time-domain surveys. Future work will extend this framework to include radial velocity cross-matching and application to other survey datasets beyond TESS.
Acknowledgements
This research was conducted entirely independently and without institutional support or funding. All computations were performed on a single personal desktop machine (Intel i3 processor, 32 GB RAM), which, despite its limitations, proved remarkably persistent. The author is deeply grateful for its service.
The author thanks the developers and maintainers of the open-source software and data archives acknowledged in the Data Availability section, which made this work possible.
Data Availability
The full CLARA pipeline, including Jupyter notebooks, source code, pre-trained URF models, and helper scripts for light curve preprocessing, anomaly scoring, and similarity matching, and the CLARA cli tool to execute the pipeline with the computational resources and efficiency as stated in section 5, is publicly available at: https://github.com/googleboy-byte/CLARA and https://github.com/googleboy-byte/clara-cli.

### v3

@@ 3.4.1 Interpretation of Cosine Similarity Results
Clustering Label TOIs Curves (%)
@@
3.4.2 Results and Interpretation of Morphological Clustering Outcomes
To explore the morphological diversity among the 620 planet-like candidates (Section 2.15), we applied a range of clustering algorithms to a 10-dimensional feature set (Section 2.15.1) engineered from phase-folded light curves. This feature set is as described in section 2.15.1. Our objective was to determine whether these features could reveal structural subgroups and isolate known TOIs within well-defined clusters. Results of clustering methods are shown in Table 5
Among all tested methods, Gaussian Mixture Models (GMM) A. P. Dempster et al. (1977) with two components produced the clearest separation: one cluster contained 17 TOIs among 155 curves, achieving the highest TOI density. Dirichlet Process GMM (DPMM) D. Görür & C. E. Rasmussen (2010) performed comparably well, grouping 16 TOIs within a smaller cluster of 114 curves, indicating a stronger enrichment but with more fragmentation overall.
While other methods such as DBSCAN, HDBSCAN, and KMeans S. Fotopoulou (2024) provided partial clustering success, they either included fewer TOIs or distributed them across broader clusters. The probabilistic nature of GMM and DPMM allowed them to better model the subtle variations in transit morphology, making them more suitable for this task. As evident from Figures 13 and 14, the planet like candidate list of 620 curves with $N-WRSS\geq 0.35$ (arbitrary threshold) filtered down to a maximum of 155 curves with morphology matching the curves of the synthetic training set used for fine tuning hyper-parameters, 17 of which are TESS Objects of Interest as listed in Table 16 and Table 17.
In conclusion, the combination of interpretable features and flexible model-based clustering yielded the best separation of high-quality transit-like signals. The GMM and DPMM results (sampled in figures 13 and 14 to visualize morphological similarity) represent the most promising configurations to guide further candidate filtering, validation, and prioritization.
@@ 3.4.2 Results and Interpretation of Morphological Clustering Outcomes
3.5 Initial Injection-Recovery Test Results
Before performing the Monte Carlo simulations of the injection-recovery test, we first injection-recovery tested each of the 10 configurations with 1980 randomly sampled real TESS curves and the 20 curves of that particular configuration for each configuration test set. The clear tradeoff trends are visible across N-WRSS thresholds in the results of this test as presented in Table 6. The $N-WRSS\geq 0.35$ achieves near-perfect recall but at the expense of $\approx 126$ candidates to vet per 20 injections, serving as an intentionally permissive baseline. Conversely, $N-WRSS\geq 0.6$ achieves an average of 84% recall with $\approx 24$ curves to vet, representing a more practical operating point. At $N-WRSS\geq 0.9$ precision is high but recall falls sharply, illustrating the cost of overly strict thresholds.
Table 6: Average recovery performance in the initial injection–recovery test. Detailed summary are present in Tables 11, 12 and 13.
N-WRSS Threshold Avg. Retrieved Avg. Total
0.35 19.6 / 20 126
0.60 16.7 / 20 24
0.90 5.3 / 20 5
3.6 Evaluation of Injection-Recovery Performance
Table 7: Monte Carlo injection–recovery performance over 80 runs at $N-WRSS\geq 0.6$. Reported values are mean $\pm$ standard deviation across runs, with 95% confidence intervals on the mean in parentheses.
Metric Mean $\pm$ SD 95% CI on Mean
Precision $0.53\pm 0.18$ [0.49, 0.57]
Recall $0.74\pm 0.24$ [0.69, 0.79]
Retrieved Injected $14.8\pm 4.8$ [13.8, 15.9]
Retrieved Total $26.1\pm 8.0$ [24.4, 27.9]
Figure 15: Comparison of precision (left) and recall (right) as functions of the detection threshold for the linear weighted sum and the normalized weighted root sum score (N-WRSS) in terms of measuring model consesus towards anomalousness. N-WRSS maintains higher recall at stricter thresholds while achieving improved precision in the high-threshold regime, indicating greater robustness against false positives.
To more rigorously quantify performance and mitigate dependence on any single injection set, we conducted over 80 Monte Carlo runs (Section 2.16.1) of the injection-recovery test using randomly sampled TESS light curves and synthetic transits. At a practical operating point of $N-WRSS\geq 0.6$, the ensemble statistics across runs (Table 7) yield a mean precision of $0.53\pm 0.18$ and mean recall of $0.74\pm 0.24$. 95% confidence intervals on these metrics were 0.490.570.49-0.57 and 0.680.790.68-0.79 respectively. On average, 14.8 $\pm$ 4.8 injected curves were successfully recovered per run, with $26.1\pm 8.0$ total candidates requiring vetting (95% confidence intervals: 13.815.913.8-15.9 recovered and 24.447.924.4-47.9 to vet). These results reinforce the earlier tradeoff trends, demonstrating that the chosen threshold balances high recall with a manageable number of candidates, while also confirming robustness across repeated randomized trials. A full visualization of performance across all thresholds is provided in Appendix 24.
To further validate the choice of N-WRSS over a simpler linear weighted sum for ensemble consensus scoring (see Section 2.11), we compared both strategies across 80+ Monte Carlo runs. Figure 15 shows mean precision and recall as functions of detection threshold. The linear weighted sum yields a sharp precision peak near 0.6, but at the cost of collapsing recall, underscoring its sensitivity to threshold choice. In contrast, N-WRSS provides smoother, more stable tradeoffs, maintaining a competitive precision while sustaining higher recall across a wider threshold range. This robustness complements its mathematical basis as a monotone concave transformation that suppresses single-model dominance in ensemble scoring.
The injection-recovery tests allow us to quantify the clustering behavior of an unsupervised probabilistic clustering method like the Gaussian Mixture Model within a Dirichlet process framework. Figure 16 and Table 8 summarize three key statistics:
  1. 1.
    Mean ratio of injected curves in the dominant cluster relative to all injected curves surviving the threshold.
  2. 2.
    The mean absolute recovery ratio. This is the number of injected curves in the dominant cluster relative to the total number of injected curves (20) regardless of N-WRSS threshold.
  3. 3.
    The mean fraction of surviving curves assigned to the dominant cluster. This is the number of curves in the dominant cluster relative to the total number of curves that survive an N-WRSS threshold.
At permissive thresholds (NWRSSN-WRSS threshold $\leq 0.35$), recovery ratios approach unity, forming small clusters (20% of curves that survive the threshold per cluster), with high recovery ratios (75-80% injected curves in this single cluster), resulting in high computational complexity. As thresholds increase ($0.5\lesssim$ N-WRSS $\lesssim 0.7$), injected curves are steered into a single dominant cluster comprising $\gtrsim 50\%$ of all surviving curves, with mean recovery ratios sustained at 60-70%. At overly strict thresholds ($N-WRSS\gtrsim 0.8$), the mean absolute recovery ratio falls below 40%, reflecting the significant loss of true signals alongside the suppression of spurious clusters. This analysis demonstrates that the URF does not simply filter anomalies but actively steers recovered injections toward coherent morphological groupings. This provides an unsupervised basis for distinguishing transit-like morphological signals from noise and empirically justifies the use of probabilistic clustering on a morphological feature set.
The false positives, as presented in Table 9, are predominantly explained by well-known astrophysical contaminants. Out of 187 non-injected curves that surpass the N-WRSS threshold and are in the dominant cluster, the largest contribution arises from binaries (100 objects), consistent with eclipsing binary light curves mimicking transit-like morphology. Generic stars (50) and variables of rotational, pulsational, or eruptive type (20) account for most of the remaining population. A small but non-negligible set of objects classified as planets or planet-like (8) demonstrates that non-injected but genuine planet-hosting curves do surpass the threshold. Finally, unresolved multiples (6) and unclassified sources (3) constitute a minor fraction. These results indicate that at the practical operating point of $N-WRSS\geq 0.6$, CLARA’s false positives are astrophysically interpretable and dominated by binaries and stellar variability.
Table 8: Quantification of probabilistic clustering on morphological feature set across NN-WRSS threshold ranges. Reported values are mean statistics across Monte Carlo runs. The Mean Ratio refers to injected curves surviving in the dominant cluster relative to all the injected curves that survive an N-WRSS threshold; the Mean Absolute Ratio is relative to the total number of injected curves (20); the Mean Fraction in Dominant Cluster is the fraction of all surviving curves assigned to the dominant cluster. The dominant cluster is the one containing the most injected curves above an N-WRSS threshold
NN-WRSS Range Mean CI Lower Mean CI Upper Mean Ratio Mean Fraction in Dominant Cluster Mean Absolute Ratio
0.35–0.40 0.51 1.00 0.83 0.24 0.76
0.40–0.50 0.42 0.98 0.74 0.37 0.74
0.50–0.65 0.37 0.86 0.62 0.51 0.68
0.65–0.80 0.37 0.97 0.59 0.51 0.54
0.80–0.95 0.30 0.97 0.57 0.45 0.32
Figure 16: Quantification of URF steering across NN-WRSS thresholds. Shown are mean recovery ratios, dominant cluster fractions, and absolute recovery ratios from Monte Carlo injection–recovery experiments. At intermediate thresholds ($0.5\lesssim N$-WRSS $\lesssim 0.7$), injected transits are preferentially steered into a single dominant cluster while maintaining $\sim$60–70% recovery, whereas stricter thresholds suppress both false positives and true signals. Numerical values are provided in Table 8.
Table 9: SIMBAD classifications of false positives (non-injected curves in dominant cluster with N-WRSS $\geq 0.6$) from Monte Carlo injection-recovery runs. Objects are grouped into broad astrophysical categories.
Category Count
Binaries 100
Variables (rot./puls./eruptive) 20
Generic Stars 50
Planets or Planet-like 8
Multiple/Unresolved systems 6
Other/Unclassified 3
Total 187
@@ 4.1 Synthesis of Key Findings
CLARA demonstrates that synthetic set design enables interpretable, tunable, and efficient URF-based anomaly detection in large astronomical surveys, linking model behavior directly to training parameters. Parameter sweeps over 36 URF-4 variants reveal transit duration as the dominant driver of TOI Recall and Anomaly Rate, and noise level as the key factor in importance-based prioritization, enabling predictive tuning within 50–300 ppm noise and 13–27 day duration ranges. The $\alpha$-parameterized scoring framework robustly controls completeness–significance trade-offs across TESS sectors, with $\alpha=0.9$ maximizing TOI recall and $\alpha=0.3$ concentrating high-value targets; combined metrics like N-WRSS unify their strengths.
@@ 4.2 Astrophysical Validation and Physical Interpretability
Anomaly–stellar parameter correlations and t-SNE embeddings show high-scoring curves form coherent astrophysical manifolds. SIMBAD cross-matching confirms that N-WRSS $\geq 0.85$ candidates always exhibit high morphological similarity to known classes, supporting physical authenticity. N-WRSS favors planet-like candidates at high thresholds, indicating appropriate sensitivity to transit morphologies.
@@ 4.3 Limitations and Scope of Applicability
Current results apply to 2-minute SPOC targets, with FFI data requiring recalibrated thresholds and parameter–performance mappings. Dependence on TOIs for validation risks bias, partly mitigated by SIMBAD-based checks and injection-recovery tests. Simplified synthetic models and truncated light curves may omit complex or long-duration anomalies, motivating more realistic generation and extended feature extraction.
@@ 4.4 Future Directions and Extensions
CLARA enables seven future research avenues:
@@
(1) scale to a 500k light curve benchmark via clara-cli to test full-pipeline scalability;
(2) implement rigorous statistical validation with confidence intervals, cross-method comparisons, and deeper analysis of why Teff, RUWE, and vtan correlate with high N-WRSS;
(3) expand morphological analysis with advanced feature combinations, deep-learned representations, and improved discovery metrics;
(4) adapt framework to other surveys (Kepler, LSST, K2) via modular synthetic generation;
(5) formalize N-WRSS theory, including square-root weighting justification, alternative fusion strategies, and a principled $\alpha$-controlled performance metric with Pareto optimization;
(6) develop automated deep learning meta-optimization for synthetic parameter tuning;
(7) integrate GAN-generated light curves for supervised post-clustering classification to filter false positives without TOI labels.
@@ 5 Conclusions
We present CLARA, a modular URF-based framework for unsupervised transit detection whose behavior is steered by synthetic training set design. Part I introduced URF-4, applied $\alpha$-dependent scoring to tune completeness ($\alpha=0.9$), significance ($\alpha=0.3$), or balance ($\alpha=0.5$), and validated consistent performance across 384k light curves from TESS Sectors 1–5. Part II used Gaia DR3 parameters, t-SNE embeddings, and SIMBAD morphological matching to confirm that N-WRSS–ranked anomalies occupy coherent astrophysical subspaces and align with known exoplanetary and stellar variability classes, with no high-score, low-similarity false positives. CLARA thus delivers a physically grounded, interpretable, and generalizable anomaly detection pipeline, extendable to other surveys and RV cross-matching.
6 Data Availability
The full CLARA pipeline, including Jupyter notebooks, source code, pre-trained URF models, and helper scripts for light curve preprocessing, anomaly scoring, and similarity matching, and the CLARA cli tool to execute the pipeline with the computational resources and efficiency as stated in section A, is publicly available at: https://github.com/googleboy-byte/CLARA and https://github.com/googleboy-byte/clara-cli.